# TrajLearn: Trajectory Prediction Learning using Deep Generative Models

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Spatial-temporal systemsccs: Computing methodologies Neural networks
@@ 1. Introduction
Figure 1. Illustrative example of the trajectory prediction problem using higher-order spatial representations (hexagons); (a) there are two potential trajectories for the pedestrian, (b) trajectories are represented on a hexagon-based tesselated map, (c) given the historical data (red) and the current location (black), two trajectories are predicted (blue and orange), (d) the actual trajectory followed.
@@
Paper Organization. Section 2 introduces preliminaries and the problem. Section 3 presents the rationale for utilizing higher-order mobility flow data, and section 4 delves into the specifics of the trajectory prediction model. Section 5 presents an empirical evaluation of our approach. Section 8 highlights ethical aspects of trajectory prediction models. Section 7 provides an overview of related work, and section 9 concludes the paper.
@@ 2.2. Problem Definition
We are now in position to formally define the problem of estimating or forecasting the future path or trajectory of an object or entity based on its current state and historical data.
@@ 3. Higher-order Mobility Flow Data
Rationale. Working with raw trajectory datasets is challenging because GPS coordinates: (i) are sparse and large amounts are needed to learn meaningful relationships, and (ii) are not very compatible (as input) with popular ML architectures, due to their continuous nature. We, therefore, propose to resort to a higher level of abstraction for representing trajectories. This transformation is done by first obtaining the routes/paths connecting raw trajectory data points through publicly available routing algorithms [^1] Open Source Routing Machine (OSRM). https://project-osrm.org/docs/v5.24.0/api/#route-service. Then, the trajectory is represented as a sequence of the higher-order elements (hexagons) traversed by the route. We favor hexagons over other rectangular partitioning methods (such as Google S2222S2 Geometry Library. http://s2geometry.io squares) because all six neighboring cells share identical properties, including equal distance to the cell’s centroid and uniform border lengths. This consistency makes hexagons more compatible with transformer architectures, as the transition from one token (hexagonal cell) to its neighbor would not be affected by any partitioning scheme. Additionally, mapping a point to its corresponding hexagonal token involves a constant-time operation through coordinate system conversions (Uber Technologies, [n. d.]).
Definition 3.1 (Map Tessellation). Let =$\mathcal{B}=\{b_{1},b_{2},...,b_{n}\}$ be a set of (regular) disjoint blocks that can fully tessellate map $\mathcal{M}$, forming a regular tiling. Each block $b_{k}\in\mathcal{B}$ is assumed to be a polygon. In our study, we opt for hexagons. A hexagon-based map tessellation offers several advantages over a grid-based one (commonly known as a tile system) (Birch et al., 2007). Note also that the tessellation can happen at different levels of resolution by defining different hexagon sizes; the smaller the hexagon size, the higher the resolution.
@@
Definition 3.2 (Higher-order Trajectory). Given a trajectory $T=p_{1}p_{2}...p_{n}$, and since every point $p_{i}$ resides within a unique block $b_{i}\in\mathcal{B}$ of a tessellated map $\mathcal{M}$, we can translate every trajectory as a sequence of blocks. The outcome is a higher-order trajectory.
By associating trajectory points with individual blocks, we imply that the predicted targets move step-wise, transitioning sequentially from one block to another. This sequence of hexagon representation enables us to establish a structured temporal framework, aligning the trajectories based on a uniform time scale.
@@
Computational Geometry. We utilize computational geometry methods to transform a map-matched trajectory into a sequence of hexagons. Every trajectory is modeled as a linestring shape type, and every hexagon as a polygon shape type. Then, their intersection can be computed using off-the-shelf methods of popular libraries. Recall that a map $\mathcal{M}$ can be tessellated using hexagons of a different size, which defines the map’s resolution. Upon data preparation, each trajectory in the dataset has been transformed into a sequence of hexagonal blocks (see Section 5.2).
@@ 4.1. Treating Trajectories as Statements
where $Q$, $K$ and $V$ are matrices representing the queries, keys and values, respectively, $W_{Q}$, $W_{K}$, and $W_{V}$ represent the respective learnable weight parameter matrices, and $d_{k}$ is the dimensionality of the keys. The matrix $\mathbf{M}$ is used to mask out future positions in the sequence, so the output of the self-attention layer for each position depends only on tokens to its left, ensuring causality in the attention mechanism. This mechanism allows the model to focus on different parts of the input sequence when generating the output. The output of the last layer is fed into layer normalization followed by a linear projection and a softmax activation that predicts the next block in the trajectory based on the probabilities of all possible next blocks:
@@
Although we followed a decoder-only transformer architecture, our approach for trajectory prediction is flexible, and other architectures used in large language models could be used. Thus, any advancements in language models are applicable and can benefit our approach with minimal effort.
@@ 4.4. Computational Complexity
Self-Attention The computational complexity of the self-attention mechanism is $O(n^{2}.d)$, where $n$ is the sequence length and $d$ is the dimension of the representation. Self-attention involves comparing each element in the sequence (like a word in a sentence) with every other element. This requires $n\times n$ comparisons or attention scores to be computed, leading to a quadratic complexity in terms of sequence length, and then, for each comparison, the model computes attention scores based on token representations, which are vectors of size $d$. This computation involves these vectors and hence introduces a factor of $d$ in the complexity.
Beam Search During the inference time, the beam search algorithm is utilized with a complexity of $O(w.\beta.k)$. Here, $w$ is the beam width, $k$ is the prediction horizon (depth of the search tree), and $\beta$ is the branching factor for each beam. At each level of the tree, the algorithm examines $w$ nodes, and for each of these nodes, it considers up to $\beta$ child nodes. However, only the best $w$ among these $w.\beta$ nodes are retained for the next level. Therefore, the number of nodes evaluated at each level is proportional to $w.\beta$, and this process repeats for $k$ levels, which is our prediction horizon.
@@ 5.1. Experimental Configuration
Computational Environment. We conducted experiments on a server equipped with an NVIDIA RTX A6000 graphics card and 320GB of memory. The model was developed in Python 3, and it was trained and deployed using the PyTorch 1.13 framework.
@@ 5.3. Baseline Methods
DeepMove (Feng et al., 2018). DeepMove is a state-of-the-art method combining a multi-modal recurrent network and a historical attention mechanism to capture both spatial and temporal dependencies. The (Wang et al., 2021c) implementation
@@
Baselines Implementation. For the baselines, we implemented the MC, LSTM, LSTM-ATTN, and GRU models ourselves. For the DeepMove model, we used the implementation provided by (Wang et al., 2021c) [^4] https://github.com/LibCity/Bigscity-LibCity, and for Flashback++ we relied on the implementation by (Deng et al., 2023) [^5] https://github.com/Pursue1221/FlashbackPlusPlus, both with adjustments to support hexagonal-based spatial data in our experiments.
@@ 5.5. Results and Discussion
Figure 7. Impact of beam width $w$ on TrajLearn’s accuracy (left) and inference time (right) on Ho-Porto, res=7.
@@
(Q5) Ablation Study. In this experiment, we investigate the impact of certain components of TrajLearn’s neural architecture. We selectively remove the beam search module and report the accuracy performance change. Table 5 shows the results, which indicate that the accuracy drops when removing the beam search. In addition, we perform a hyperparameter analysis to gauge the impact of various parameters on accuracy: embedding dimension, number of decoder layers, and number of attention heads. These tests were carried out using the GeoLife@7 dataset. Figure 9 presents the results. From our observations, TrajLearn’s performance benefits from an increase in the number of layers, as it allows for a more comprehensive processing of dependencies. Likewise, enhancing the number of attention heads results in a better outcome. Additionally, increasing the embedding dimension enhances performance up to a certain threshold, likely due to the constrained input length.
@@ 5.6. Interpretability Study
Based on Figure 11, a few observations can be made: 1) some positions, notably hexagon 4, play a crucial role across multiple heads, suggesting they contain important information for the trajectory sequence. 2) There is evidence of significant attention between non-adjacent positions, indicating that the model identified some long-distance relationships. For instance, as depicted in Figure 11(a), hexagon 4 represents a turn on the highway that significantly could influence the prediction of the next block (11) based on recent hexagons. Moreover, we have used the max function to aggregate the attention weights, highlighting the highest attention score between each pair of positions across all heads in Figure 11(b).

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TSASisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Spatial-temporal systemsccs: Computing methodologies Neural networks
@@ 1. Introduction
Figure 1. Illustrative example of the trajectory prediction problem using higher-order spatial representations (hexagons); (a) there are two potential trajectories for the pedestrian, (b) trajectories are represented on a hexagon-based tessellated map, (c) given the historical data (red) and the current location (black), two trajectories are predicted (blue and orange), (d) the actual trajectory followed.
[…]
A multi-panel figure illustrating trajectory prediction using a hexagon-based tessellated map. In panel (a), a pedestrian faces two potential trajectories. Panel (b) shows these trajectories mapped onto a grid of hexagons. In panel (c), historical data (in red) and the current position (in black) are used to predict two future trajectories, highlighted in blue and orange. Panel (d) displays the actual path followed.
@@
Paper Organization. Section 2 introduces preliminaries and the problem. Section 3 presents the rationale for utilizing higher-order mobility flow data, and section 4 delves into the specifics of the trajectory prediction model. Section 5 presents an empirical evaluation of our approach. Section 6 introduces the use of hierarchical maps for trajectory prediction in complex urban environments. Section 7 provides an overview of related work. Section 8 highlights ethical aspects of trajectory prediction models, and section 9 concludes the paper.
@@ 2.2. Problem Definition
We are now in a position to formally define the problem of estimating or forecasting the future path or trajectory of an object or entity based on its current state and historical data.
@@ 3. Higher-order Mobility Flow Data
Rationale. Working with raw trajectory datasets is challenging because GPS coordinates: (i) are sparse, and large amounts are needed to learn meaningful relationships, and (ii) are not very compatible (as input) with popular ML architectures, due to their continuous nature. We, therefore, propose to resort to a higher level of abstraction for representing trajectories. This transformation is done by first obtaining the routes/paths connecting raw trajectory data points through publicly available routing algorithms [^1] Open Source Routing Machine (OSRM). https://project-osrm.org/docs/v5.24.0/api/#route-service. Then, the trajectory is represented as a sequence of the higher-order elements (hexagons) traversed by the route. We favor hexagons over other rectangular partitioning methods (such as Google S2222S2 Geometry Library. http://s2geometry.io squares) because all six neighboring cells share identical properties, including equal distance to the cell’s centroid and uniform border lengths. Moreover, hexagon-based tessellations offer several additional advantages: They provide a simpler and more symmetric definition of the nearest neighborhood, as each hexagon has six equidistant neighbors, eliminating the ambiguity inherent in rectangular grids that possess two types of neighbors (orthogonal and diagonal) with differing distances. Hexagons also approximate circles more closely than squares, resulting in a lower perimeter-to-area ratio (for instance, a unit-area hexagon has a perimeter of approximately 3.722 compared to 4 for a square), which minimizes edge effects. Furthermore, hexagonal grids exhibit greater isotropy, meaning that the relationship between grid-based and Euclidean distances varies less with direction, thereby reducing bias in spatial measurements and modeling dispersal and connectivity. This consistency makes hexagons more compatible with transformer architectures, as the transition from one token (hexagonal cell) to its neighbor would not be affected by any partitioning scheme. Additionally, mapping a point to its corresponding hexagonal token involves a constant-time operation through coordinate system conversions (Uber Technologies, [n. d.]).
Definition 3.1 (Map Tessellation). Let =$\mathcal{B}=\{b_{1},b_{2},...,b_{n}\}$ be a set of (regular) disjoint blocks that can fully tessellate the map $\mathcal{M}$, forming a regular tiling. Each block $b_{k}\in\mathcal{B}$ is assumed to be a polygon. In our study, we opt for hexagons. A hexagon-based map tessellation offers several advantages over a grid-based one (commonly known as a tile system) (Birch et al., 2007). Note also that the tessellation can happen at different levels of resolution by defining different hexagon sizes; the smaller the hexagon size, the higher the resolution.
@@
[…]
A schematic diagram illustrating the process of constructing higher-order trajectory data. The figure depicts the transformation of raw trajectory inputs into enhanced representations.
Definition 3.2 (Higher-order Trajectory). Given a trajectory $T=p_{1}p_{2}...p_{n}$, and since every point $p_{i}$ resides within a unique block $b_{i}\in\mathcal{B}$ of a tessellated map $\mathcal{M}$, we can translate every trajectory as a sequence of blocks. By associating trajectory points with individual blocks, we imply that the predicted targets move step-wise, transitioning sequentially from one block to another. The outcome is a higher-order trajectory.
@@
Computational Geometry. We utilize computational geometry methods to transform a map-matched trajectory into a sequence of hexagons. Every trajectory is modeled as a linestring shape type, and every hexagon as a polygon shape type. Then, their intersection can be computed using off-the-shelf methods of popular libraries. Recall that a map $\mathcal{M}$ can be tessellated using hexagons of different sizes, which defines the map’s resolution. Upon data preparation, each trajectory in the dataset has been transformed into a sequence of hexagonal blocks (see Section 5.2).
@@ 4. Trajectory Prediction Learning
[…]
A schematic diagram illustrating the high-level architecture of the proposed framework.
@@ 4.1. Treating Trajectories as Statements
where $Q$, $K$ and $V$ are matrices representing the queries, keys and values, respectively, $W_{Q}$, $W_{K}$, and $W_{V}$ represent the respective learnable weight parameter matrices, and $d_{k}$ is the dimensionality of the keys. The matrix $\mathbf{M}$ is used to mask out future positions in the sequence, so the output of the self-attention layer for each position depends only on tokens to its left, ensuring causality in the attention mechanism. This mechanism allows the model to focus on different parts of the input sequence when generating the output. The output of the last layer is fed into layer normalization, followed by a linear projection and a softmax activation that predicts the next block in the trajectory based on the probabilities of all possible next blocks:
@@
Although we followed a decoder-only transformer architecture, our choice is motivated by the autoregressive nature of trajectory prediction, which requires the sequential generation of future spatial tokens conditioned solely on past trajectory data. A decoder-only transformer directly models the conditional distribution of each future hexagonal block given the preceding sequence, thereby simplifying both training and inference. This architecture also reduces computational overhead and enables seamless integration with our constrained beam search mechanism (subsection 4.3) to enforce spatial continuity between adjacent blocks. While encoder-decoder architectures may be beneficial for tasks involving more complex input-output mappings, our focus on the trajectory prediction task, renders the decoder-only approach particularly efficient. Nonetheless, our approach for trajectory prediction is flexible, and other architectures used in large language models could be used. Thus, any advancements in language models are applicable and can benefit our approach with minimal effort.
@@ 4.3. Beam Search with Constraints
[…]
Comparison of training with and without teacher forcing. The top diagram shows a model trained without teacher forcing, where predictions are fed back into the model. The bottom diagram shows training with teacher forcing, where ground truth is used at each step instead of predicted outputs.
@@
[…]
A beam search example with beam width 2 and sequence length 3. The figure illustrates multiple branching paths, showing how the top-k sequences are maintained at each step of the search process.
@@ 4.4. Computational Complexity
Self-Attention. The computational complexity of the self-attention mechanism is $O(n^{2}.d)$, where $n$ is the sequence length and $d$ is the dimension of the representation. Self-attention involves comparing each element in the sequence (like a word in a sentence) with every other element. This requires $n\times n$ comparisons or attention scores to be computed, leading to a quadratic complexity in terms of sequence length, and then, for each comparison, the model computes attention scores based on token representations, which are vectors of size $d$. This computation involves these vectors and hence introduces a factor of $d$ in the complexity.
Beam Search. During the inference time, the beam search algorithm is utilized with a complexity of $O(w.\beta.k)$. Here, $w$ is the beam width, $k$ is the prediction horizon (depth of the search tree), and $\beta$ is the branching factor for each beam. At each level of the tree, the algorithm examines $w$ nodes, and for each of these nodes, it considers up to $\beta$ child nodes. However, only the best $w$ among these $w.\beta$ nodes are retained for the next level. Therefore, the number of nodes evaluated at each level is proportional to $w.\beta$, and this process repeats for $k$ levels, which is our prediction horizon.
Hexagonal Tessellation. Mapping a GPS coordinate to its corresponding hexagon is performed in constant time, i.e., $O(1)$, due to efficient coordinate-to-hexagon conversion methods. This ensures that the transformation from raw trajectory data to a hexagonal representation introduces minimal computational overhead. While the map-matching process, which aligns raw GPS data with the appropriate road network or spatial features, can be computationally intensive when processing large-scale data for training and testing, it is executed as part of an offline preprocessing pipeline. This separation guarantees that these heavy-lifting tasks do not impact the speed of the online inference process. Moreover, for real-time trajectory inference in practical applications, incremental map-matching techniques can be employed to update the trajectory representation dynamically as new data arrives, without significant overhead. Consequently, although the offline preprocessing stage may be resource-intensive, it is decoupled from real-time operations, ensuring that TrajLearn remains well-suited for online or real-time applications.
@@ 5. Experimental Evaluation
  5. (Q5)
    Ablation Study. How does beam search with constraint and model architecture hyperparameters impact the model’s performance?
@@ 5.1. Experimental Configuration
Computational Environment. We conducted experiments on a server equipped with an NVIDIA RTX A6000 graphics card and 320GB of memory. The model was developed in Python 3 and trained and deployed using the PyTorch 1.13 framework.
@@ 5.3. Baseline Methods
DeepMove (Feng et al., 2018). DeepMove is a state-of-the-art method combining a multi-modal recurrent network and a historical attention mechanism to capture both spatial and temporal dependencies.
@@
Baselines Implementation. For the baselines, we implemented the MC, LSTM, LSTM-ATTN, and GRU models ourselves. For the DeepMove model, we used the implementation provided by (Wang et al., 2021c) [^4] https://github.com/LibCity/Bigscity-LibCity, and for Flashback++ we relied on the implementation by (Deng et al., 2023) [^5] https://github.com/Pursue1221/FlashbackPlusPlus, both with adjustments to support hexagonal-based spatial data in our experiments. Computational complexity, memory complexity, and the number of trainable parameters comparison of baseline models and TrajLearn is reported in Table 5.
@@
Model Time Complexity Memory Complexity # Parameters
MC $O(1)$ $O(1)$ $\approx 10K$
LSTM $O(T\cdot H^{2})$ $O(T\cdot H)$ $\approx 6.1M$
LSTM-ATTN $O(T\cdot H^{2}+T^{2}\cdot H)$ $O(T^{2}\cdot H)$ $\approx 6.1M$
GRU $O(T\cdot H^{2})$ $O(T\cdot H)$ $\approx 5.1M$
DeepMove $O(T\cdot H^{2}+T^{2}\cdot H)$ $O(T^{2}\cdot H)$ $\approx 8.4M$
Flashback++ $O(T\cdot H^{2})$ $O(T\cdot H)$ $\approx 6.5M$
TrajLearn $O(T\cdot H^{2}+T^{2}\cdot H)$ $O(T^{2}\cdot H)$ $\approx 7.3M$
Table 5. Computational complexity, memory complexity, and number of trainable parameters of baseline models and TrajLearn for a single inference. Here, $T$ is the sequence length, and $H$ is the hidden size. The trainable parameter count is reported for models trained on Ho-GeoLife, res=7.
@@ 5.5. Results and Discussion
[…]
A heatmap visualization divided into three panels displaying the performance of the TrajLearn model on the Ho-Porto dataset (resolution 7). Panel (a) shows Accuracy@1, panel (b) shows Accuracy@3, and panel (c) shows Accuracy@5. The horizontal axis represents the prediction horizon (k) while the vertical axis represents the input length (l). Color intensity in each panel indicates the accuracy achieved by the model.
@@
Figure 7. Impact of beam width $w$ on TrajLearn’s accuracy (left) and inference time (right) on Ho-Porto, res=7 and batch size of 64
[…]
@@
[…]
Six-panel figure arranged in two columns and three rows. Each row corresponds to a different dataset: the top row shows results for Ho-Porto, the middle row for Ho-Rome, and the bottom row for Ho-GeoLife. In each row, the left panel plots Accuracy@1 as a function of the prediction horizon, while the right panel plots Accuracy@1 as a function of the actual distance traveled. Different markers and colors denote the resolution levels: red star for resolution 7, blue bullet for resolution 8, and black triangle for resolution 9.
(Q5) Ablation Study. In this experiment, we investigate the impact of certain components of TrajLearn’s neural architecture. We selectively remove the beam search module and report the accuracy performance change. Table 6 shows the results, which indicate that the accuracy drops when removing the beam search. In addition, we perform a hyperparameter analysis to gauge the impact of various parameters on accuracy: embedding dimension, number of decoder layers, and number of attention heads. These tests were carried out using the Ho-GeoLife dataset at resolution 7. Figure 9 presents the results. From our observations, TrajLearn’s performance benefits from an increase in the number of layers, as it allows for a more comprehensive processing of dependencies. Likewise, enhancing the number of attention heads results in a better outcome. Additionally, increasing the embedding dimension enhances performance up to a certain threshold, likely due to the constrained input length.
@@
[…]
A three-column figure composed of nine panels arranged in three rows. The top row shows results for the Ho-GeoLife dataset, the middle row for Ho-Rome, and the bottom row for Ho-Porto, all at resolution 7. The left column displays plots of Accuracy@1 with varying embedding vector sizes, the middle column shows plots for different numbers of attention heads, and the right column presents plots for varying numbers of Transformer layers.
@@
Discussion. Beyond addressing the five experimental evaluation questions, we now delve deeper into the underlying reasons for TrajLearn’s performance improvements. Unlike conventional autoregressive approaches, such as RNN-based models (e.g., LSTM and GRU), which often suffer from vanishing gradients and limited long-range context, TrajLearn leverages a transformer-based architecture that models the joint distribution of future trajectories, thereby effectively capturing complex higher-order mobility flows. The inherent self-attention mechanism of transformers enables the model to learn intricate spatial-temporal dependencies over extended sequences, seamlessly integrating cues from both local (recent) contexts and long-range dependencies that are critical for accurate prediction, thus yielding more informed and precise outcomes. Furthermore, the model’s generative formulation facilitates the exploration of multiple plausible future paths in a coherent and probabilistic manner, enhancing its ability to account for the inherent uncertainty and multimodality in movement data. Moreover, although hexagonal tessellation is common across all evaluated models, its superior ability to capture subtle spatial dependencies—thanks to its uniform neighborhood structure—provides a richer and more consistent representation of the movement space. This enhanced spatial representation is particularly well-exploited by a powerful model like the transformer, which can discern nuanced relationships both between adjacent spatial units and across larger regions. Additionally, the employment of teacher forcing during training guides the model with ground-truth sequences, thereby promoting more robust learning of complex spatial-temporal patterns. Finally, the integration of a constrained beam search mechanism during inference enables TrajLearn to explore multiple plausible future paths in a coherent and probabilistic manner, ensuring that predicted trajectories adhere to spatial continuity constraints. This approach effectively accounts for the inherent uncertainty and multimodality in movement data, ultimately generating realistic and feasible paths. Collectively, these design choices empower TrajLearn to handle complex higher-order mobility flows, leading to performance improvements of up to approximately 40% over baseline methods, as demonstrated in our experiments.
@@ 5.6. Interpretability Study
[…]
A composite image displaying heatmaps for the attention weights of all 8 heads when predicting hexagon 11. Each heatmap visualizes how a specific attention head distributes its focus across input elements, illustrating the internal workings of the model during this prediction step.
@@
[…]
A two-panel figure demonstrating how TrajLearn predicts a future trajectory. Panel (a) shows an example trajectory where the input sequence consists of hexagons 1 through 10, and the model predicts hexagon 11 as the next step. Panel (b) presents an aggregated heatmap of the attention weights from all 8 heads, illustrating how the model focuses on different parts of the input during the prediction process.
@@
Based on Figure 11(b), a few observations can be made: 1) some positions, notably hexagon 4, play a crucial role across multiple heads, suggesting they contain important information for the trajectory sequence. 2) There is evidence of significant attention between non-adjacent positions, indicating that the model identified some long-distance relationships. For instance, as depicted in Figure 11(a), hexagon 4 represents a turn on the highway that significantly could influence the prediction of the next block (11) based on recent hexagons. Moreover, we have used the max function to aggregate the attention weights, highlighting the highest attention score between each pair of positions across all heads in Figure 11(b).